# SaVe-Tag: Semantic-aware Vicinal Risk Minimization for Long-Tailed Text-Attributed Graphs

### v3

@@ 4.1.2 Manifold-preserving Interpolation
Proof.
By conditioning the model on the target class $c$, prompting transforms $\mathcal{L}_{\tau}$ into a class-conditional generator: $\hat{t}\sim p\bigl{(}t\mid t_{1},t_{2},c\bigr{)}.$ If the model has learned the conditional support of $\mathcal{M}_{c}$ accurately—e.g., through ample in-class examples—then with probability at least $1-\delta$, the sampled $\hat{t}$ will lie within the semantic manifold of class $c$, i.e. $\phi(\hat{t})\in\mathcal{M}_{c}$. ∎
4.1.3 Boundary-enriching Effect of LLM Samples
Intuitively, the greatest gain from augmentation arises when synthetic points fill sparsely-populated decision-boundary zones [15]. We therefore quantify how many generated texts land in these border regions and later show (Theorem 4.6) that a higher fraction yields a lower vicinal risk.
Let $\gamma(x)=f_{\theta}^{(y)}(x)-\max_{j\neq y}f_{\theta}^{(j)}(x)$ be the logit margin; write $\gamma_{\min}(S)=\min_{x\in S}\gamma(x)$ for a set $S$ [16]. Denote the augmented training set by $\tilde{\mathcal{D}}=\{(x_{i},y_{i})\}_{i=1}^{n}\cup\{(\hat{x}_{j},\hat{y}_{j})\}_{j=1}^{m}.$
Definition 4.3 (Boundary-Coverage Rate (BCR)).
Following [17], a synthetic example $\hat{x}$ is a boundary sample (a borderline) iff the $k$-NN majority label on $\mathcal{M}$ differs from $\hat{y}$. BCR is the fraction of such samples in $\tilde{\mathcal{D}}$.
Theorem 4.4 (Margin lower bound).
Fix $\delta>0$. Let $f_{\theta}^{\mathrm{orig}}$ be trained on $\mathcal{D}$ with margin $\gamma_{\min}(\mathcal{D})=\gamma_{0}$. Augment with $m$ LLM-generated points and retrain to obtain $f_{\theta}^{\mathrm{aug}}$. Assume each boundary sample satisfies $\lvert\gamma(\hat{x}_{j})\rvert\leq\delta$. Then, adapting the proof of [18],
$\gamma_{\min}\!\bigl{(}\tilde{\mathcal{D}}\bigr{)}\;\;\geq\;\;\gamma_{0}-\delta\bigl{(}1-\mathrm{BCR}\bigr{)}.$
Hence more boundary samples $\Rightarrow$ a tighter realized margin.
Next, we show that such boundary-rich distribution can also minimize vicinal risk.
4.1.4 Vicinal Risk Minimization with an LLM Neighborhood
Vicinal Risk Minimization (VRM) defines a local surrogate loss around each training point. When the neighborhood is populated with boundary samples, this surrogate becomes a sharper proxy for the true generalization error. We formalize the link below.
Theorem 4.5 (On-manifold Vicinal Risk Theorem).
Let $\{(t_{i},c_{i})\}_{i=1}^{n}$ be the training set and define the LLM-based vicinal distribution
$\mathcal{V}_{\mathcal{L}_{\tau}}(t_{i})=\sum_{j=1}^{n}\alpha_{ij}\,\delta_{\,\hat{t}=\mathcal{L}_{\tau}(t_{i},t_{j}\mid c_{i})},\qquad\sum_{j}\alpha_{ij}=1.$
The corresponding vicinal risk is $R_{vrm}(f_{\theta})=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\hat{t}\sim\mathcal{V}_{\mathcal{L}_{\tau}}(t_{i})}\!\bigl{[}\mathbf{L}\bigl{(}f_{\theta}(\phi(\hat{t})),c_{i}\bigr{)}\bigr{]}.$
If  Theorem 4.2 holds, $R_{vrm}(f_{\theta})$ is evaluated on-manifold with probability $\geq 1-\delta$.
Theorem 4.6 (Boundary-coverage $\Rightarrow$ vicinal-risk reduction).
Let $f_{\theta}^{\mathrm{aug}}$ be obtained from VRM training on $\tilde{\mathcal{D}}$. Denote by $L$ the Lipschitz constant of loss $\mathbf{L}$ in its first argument. Then, for any $\eta>0$,
$R_{vrm}(f_{\theta}^{\mathrm{aug}})\;\;\leq\;\;R_{vrm}(f_{\theta}^{\mathrm{orig}})\;-\;L\,\gamma_{0}\,\bigl{(}\mathrm{BCR}-\eta\bigr{)}\;+\;\mathcal{O}(\delta).$
In words: increasing BCR lowers the vicinal risk up to constants dictated by the original margin and the loss smoothness. Hence the boundary-enriching effect of LLM samples (Sec. 4.1.3) directly translates to a tighter VRM bound. See Appendix F for detailed proofs.
4.2 Graph as Natural Filters
We now show that our edge assignment mechanism with confidence function naturally acts as a filter: well-aligned synthetic nodes are pulled toward the correct class manifold, whereas noisy nodes are automatically isolated. Throughout this section we retain the notation introduced in Section 4.1.1.
4.2.1 Setup and Notation
Let $\bar{\mathcal{V}}$ be the set of synthetic nodes with textual attributes $\bar{\mathcal{T}}$ after interpolation. Define $\mathcal{V}^{\prime}=\mathcal{V}\cup\bar{\mathcal{V}},~{}\mathcal{T}^{\prime}=\mathcal{T}\cup\bar{\mathcal{T}}.$ We then form the augmented graph $\mathcal{G}^{\prime}=\bigl{(}\mathcal{V}^{\prime},\,\mathcal{E}^{\prime},\,\mathcal{T}^{\prime}\bigr{)},$ where $\mathcal{E}^{\prime}\setminus\mathcal{E}$ is obtained by evaluating, for each synthetic node $\hat{v}\in\bar{\mathcal{V}}$ and each node $u\in\mathcal{V}^{\prime}$, whether to include the edge $(\hat{v},u)$.
Aggregator.
Following [19, 20], any message-passing GNN layer is abstracted as
$h^{(\ell+1)}_{v}\;=\;\alpha\,\mathbf{W}\,h^{(\ell)}_{v}\;+\;(1-\alpha)\!\!\sum_{u\in\mathcal{N}(v)}\!\beta_{vu}\,\mathbf{W}\,h^{(\ell)}_{u},$ (1)
where $h^{(\ell)}_{v}\!\in\!\mathbb{R}^{p}$, $\alpha\!\in\![0,1]$, $\beta_{vu}\!\geq\!0$ and $\sum_{u\in\mathcal{N}(v)}\beta_{vu}=1$.
@@ 4.2.1 Setup and Notation
After training, each class $c$ occupies a roughly low-dimensional manifold $\mathcal{M}_{c}\!\subset\!\mathbb{R}^{p}$ in representation space [21]. Following [22], we write $\text{dist}(x,\mathcal{S}):=\inf_{y\in\mathcal{S}}\|x-y\|_{2}$.
4.2.2 Confidence–based Edge Function
Definition 4.7 (Confidence function).
A measurable map $\kappa:\mathbb{R}^{d^{\prime}}\!\times\!\mathbb{R}^{d^{\prime}}\!\to\![0,1]$ assigns a confidence score to any pair of text embeddings. Given a threshold $\tau\!\in\!(0,1)$ (or a global top–$k$ rule), we add $(u,v)$ to $\mathcal{E}^{\prime}$ iff $\kappa(z_{u},z_{v})\geq\tau$.
4.2.3 Confidence $\Rightarrow$ Pulling
Theorem 4.8 (Pulling well-aligned nodes).
Let $\hat{v}$ be a synthetic node labelled $c$ and it has a nonempty set of neighbors $\mathcal{N}^{\prime}(\hat{v})=\{u\in V:\kappa(z_{\hat{v}},z_{u})\geq\tau\}\neq\varnothing$. Assume every neighbor $u\in\mathcal{N}^{\prime}(\hat{v})$ satisfies $\text{\emph{dist}}\bigl{(}h^{(\ell)}_{u},\mathcal{M}_{c}\bigr{)}\leq\varepsilon$. Based on  [23, 24], there exists $\gamma\!\in\!(0,1)$, independent of $\kappa$, such that
$\text{\emph{dist}}\bigl{(}h^{(\ell+1)}_{\hat{v}},\mathcal{M}_{c}\bigr{)}\;\leq\;\gamma\;\text{\emph{dist}}\bigl{(}h^{(\ell)}_{\hat{v}},\mathcal{M}_{c}\bigr{)}.$
Iterating the Equation (1) yields exponential convergence of $h^{(\ell)}_{\hat{v}}$ toward $\mathcal{M}_{c}$  [25].
4.2.4 No confidence $\Rightarrow$ Isolation
Theorem 4.9 (Isolating noisy nodes).
If $\kappa(z_{\hat{v}},z_{u})<\tau$ for every $u\!\in\!V$ then $\mathcal{N}^{\prime}(\hat{v})=\varnothing$ and $h^{(\ell+1)}_{\hat{v}}=\alpha\,\mathbf{W}\,h^{(\ell)}_{\hat{v}}\,$ for all layers $\ell$. Consequently, $\hat{v}$ neither influences nor is influenced by $\mathcal{G}^{\prime}$.
Proof.
Immediately from Definition 4.7 and Equation 1. ∎
5 Empirical Results
In this section, we present extensive experiments to validate our proposed method, emphasizing its benefits from both Semantic and Structural perspectives. We further evaluate the integrated framework to demonstrate its effectiveness in addressing node classification on long-tailed graphs. Specifically, our experiments aim to answer the following three research questions (RQs):
@@ 5 Empirical Results
5.1 Experimental Setup
5.1.1 Datasets
@@
Following prior studies on long-tailed graph learning [26, 27, 9], we use Cora [28], PubMed [29], and Citeseer [29], along with three larger Amazon datasets, namely Photo, Computer, and Children [30], for scalability analysis.
@@
5.1.2 Metrics
Following prior work on long-tailed graph classification [3, 2, 26, 27, 9, 8], we evaluate model performance using four standard metrics: balanced accuracy (bAcc), macro-F1 score (F1), geometric mean (GMean), and traditional accuracy (Acc). All experiments are repeated five times with different fixed seeds, and we report the mean performance across these runs.
To further assess boundary-enriching effects, we report the Boundary Coverage Ratio (BCR), as defined in Definition 4.3, and the Boundary Proximity Score (BPS). BPS is computed as the reciprocal of the Trust Score [31], measuring the ratio between the distance to the nearest out-of-class centroid and the distance to the in-class centroid. Additionally, we report the In-Class Rate (ICR), defined as the percentage of samples predicted with the same label as their ground truth. Implementation details for these metrics are provided in Section 5.2.
5.1.3 Baselines
We compare our method against two categories of baselines: (1) LLM-based data augmentation, including zero shot and few shot approaches [10, 11]; and (2) long-tailed graph learning, including Oversampling [32], SMOTE [7], Embed-SMOTE[33], MixupForGraph [9], GraphSMOTET[8], GraphSMOTEO[8], LTE4G[3], and HierTail [2]. See Appendix B for details.
To comprehensively evaluate our method, we design and implement three variants of SaVe-Tag, each corresponding to a specific VRM-based strategy. Detailed prompting strategies for each variant are provided in the Appendix B.
  • SaVe-TagS: Inspired by SMOTE [7], interpolates with nearest neighbors within the same class.
  • SaVe-TagM: Extends SMOTE [7] by interpolating embeddings with nearest neighbors from potentially different classes, inspired by Mixup techniques.
5.1.4 Implementation Details
5.2 Semantic-aware Benefits (RQ1)
On-Manifold. To compare on-manifold LLM-based and off-manifold numeric interpolation, we evaluate node classification on multiple benchmarks. Figure 4(a) shows results on Citeseer. Embed leverages LLaMA 3.2-1B embeddings for interpolation, while SimCSE and SBERT serve as alternative embedding-based baselines. Our LLM-based method using LLaMA 3.2-1B consistently outperforms all numeric interpolation baselines. Similar trends on other datasets can be found in Appendix A.
Figure 4: LLM interpolation preserves manifolds and enriches boundaries on the Citeseer dataset.
Method Cora Pubmed Citeseer Acc F1 GMean bAcc Acc F1 GMean bAcc Acc F1 GMean bAcc Original 72.5$\pm$1.4 70.4$\pm$1.1 84.5$\pm$0.6 74.7$\pm$0.9 55.1$\pm$8.2 47.8$\pm$15.0 62.2$\pm$7.9 51.5$\pm$9.7 53.7$\pm$0.1 49.6$\pm$2.2 71.4$\pm$0.9 55.9$\pm$1.3 Zero-shot 68.3$\pm$0.5 63.3$\pm$0.7 81.1$\pm$0.4 69.4$\pm$0.7 71.4$\pm$0.4 71.4$\pm$0.5 78.3$\pm$0.4 72.3$\pm$0.5 64.8$\pm$0.5 62.1$\pm$0.5 78.0$\pm$0.3 65.2$\pm$0.5 Few-shots 73.9$\pm$0.6 72.7$\pm$0.6 85.4$\pm$0.2 76.1$\pm$0.3 72.0$\pm$0.3 72.0$\pm$0.3 79.2$\pm$0.2 73.6$\pm$0.3 68.2$\pm$0.9 65.6$\pm$0.7 79.8$\pm$0.4 67.8$\pm$0.5 LLMO 73.7$\pm$0.7 72.1$\pm$0.7 85.3$\pm$0.3 76.1$\pm$0.4 71.5$\pm$0.3 71.4$\pm$0.3 78.8$\pm$0.3 73.1$\pm$0.4 69.6$\pm$0.8 67.0$\pm$0.6 80.5$\pm$0.4 68.8$\pm$0.5 LLMM 74.2$\pm$0.4 72.6$\pm$0.3 85.6$\pm$0.1 76.5$\pm$0.2 71.9$\pm$0.4 72.0$\pm$0.4 79.0$\pm$0.2 73.4$\pm$0.3 69.5$\pm$0.6 66.8$\pm$0.4 80.3$\pm$0.2 68.6$\pm$0.3 LLMS 74.5$\pm$0.4 72.5$\pm$0.5 85.5$\pm$0.2 76.3$\pm$0.3 72.5$\pm$0.4 72.5$\pm$0.4 79.4$\pm$0.3 73.8$\pm$0.3 69.2$\pm$0.6 66.5$\pm$0.5 80.1$\pm$0.3 68.3$\pm$0.5
Table 1: Comparison of VRM-based LLM interpolation with standard LLM augmentation. Subscripts O, M, and S denote the prompting strategies of their corresponding SaVe-Tag variants, but without edge assignment. Best and second-best results are emboldened and underlined, respectively.
Boundary-enriching Effects. We first visualize the embeddings from LLM-based and numeric interpolation. Figure 4(b) shows the UMAP projection of the Citeseer dataset, with the boundary of the original data circled. While numeric interpolated samples tend to stay within the original boundary, LLM-generated samples often extend beyond it, enlarging the VRM decision region.
We continue the evaluation with BCR and BPS metrics, as shown in the Figure 4c and the Appendix A. These metrics are computed for three settings: the original data, the combined set of original and interpolated data (Interp+Orig), and the interpolated data alone (Interp-Only). The results indicate that the original data typically achieves the highest BCR and BPS values, suggesting that real-world data tends to lie near decision boundaries. Notably, our LLM-based interpolation consistently outperforms its numerically interpolated counterpart in both metrics, better aligned with the original distribution with enriched boundary regions. .
In addition, we train a separate MLP classifier for each dataset to evaluate the class consistency of the augmented data. Each classifier is trained on a balanced dataset, where all classes have an equal number of training samples. Figure 4(d) shows the predictions on augmented samples from Citeseer, where the classifier achieves an accuracy of $71.3\pm 0.7$. As shown, numerically interpolated samples achieve 100% In-Class Rate (ICR) for all augmented classes, indicating complete class consistency. In contrast, LLM-interpolated samples yield ICRs of 81%, 69%, 94%, and 100%, suggesting that some synthesized samples lie near class boundaries and may be predicted as neighboring classes. Additional results on other datasets are provided in the Appendix A.
LLM as Vicinal Risk Minimization. We evaluate the effectiveness of LLM-based interpolation within the Vicinal Risk Minimization (VRM) framework. As shown in Table 1, interpolation-based prompting significantly enhances the quality of synthesized textual data, resulting in improved performance on downstream node classification tasks.
5.3 Necessity of Confidence-based Edge Assignment (RQ2)

### v5

@@ 3.3. Graph as a Natural Filter
Global top-$K$ selection. Construct the score matrix over all pairs $\mathcal{P}:=\mathcal{V}^{\prime}\times\bar{\mathcal{V}}$ and select the global top-$K$:
[…]
Here $K=k\,|\bar{\mathcal{V}}|$, yielding an average of $k$ edges per synthetic node. The final augmented graph is $\mathcal{G}^{\prime}\;:=\;(\mathcal{V}^{\prime},\,\mathcal{E}\cup\bar{\mathcal{E}}),$ which is used for downstream node classification.
Interpretation. Why confidence?
Noisy generations exist. The learned gate $\kappa$ (calibrated on real edges) prioritizes reliable anchors: high-agreement neighborhoods pull a synthetic node toward the correct region, while uniformly low agreement leaves it isolated, preventing error propagation.
3.3.3. Denoising in GNN Aggregation.
Aggregator. Adapting (Kipf and Welling, 2017; veličković2018graphattentionnetworks), consider a generic message-passing layer that updates the hidden representation $h^{(\ell)}_{v}\in\mathbb{R}^{p}$ of node $v$ at layer $\ell$ by blending its own transformed features with a weighted sum of its neighbors’ transformed features:
(2) $h^{(\ell+1)}_{v}\;=\;\alpha\,\mathbf{W}\,h^{(\ell)}_{v}\;+\;(1-\alpha)\!\!\sum_{u\in\mathcal{N}(v)}\!\beta_{vu}\,\mathbf{W}\,h^{(\ell)}_{u}$,
where $\alpha\in[0,1]$ is a mixing coefficient, $\mathbf{W}\in\mathbb{R}^{p\times p}$ is a trainable weight matrix, $\mathcal{N}(v)$ denotes the neighbor set of $v$ in $\mathcal{G}^{\prime}$, and $\beta_{vu}\geq 0$ are neighbor weights with $\sum_{u\in\mathcal{N}(v)}\beta_{vu}=1$.
@@
After training, each class $c\in\{1,\dots,C\}$ forms a representation manifold $\mathcal{M}_{c}^{\prime}\subset\mathbb{R}^{p}$ in the learned feature space. For any set $\mathcal{S}\subset\mathbb{R}^{p}$, write $\mathrm{dist}(\xi,\mathcal{S})=\inf_{y\in\mathcal{S}}\|\xi-y\|_{2}$ for the Euclidean distance from a point $\xi$ to the closest point in $\mathcal{S}$.
Theorem 3.8 ( Confidence $\Rightarrow$ pulling ).
Let $\hat{v}\in\bar{\mathcal{V}}$ be synthetic with label $c$ and define the confidence-filtered anchor set (for threshold $\tau\in(0,1)$)
$\mathcal{N}(\hat{v})\;:=\;\{\,u\in\mathcal{V}:\kappa(\hat{v},u)\geq\tau\,\}\neq\varnothing.$
[…]
$\mathrm{dist}\bigl(h^{(\ell+1)}_{\hat{v}},\mathcal{M}_{c}^{\prime}\bigr)\;\leq\;\varrho\;\mathrm{dist}\bigl(h^{(\ell)}_{\hat{v}},\mathcal{M}_{c}^{\prime}\bigr),$
and iterating (2) yields exponential convergence toward $\mathcal{M}_{c}^{\prime}$ under standard non-expansiveness conditions on (2) (Li et al., 2018; Gama et al., 2020; Cordonnier et al., 2025).
Interpretation. High-confidence anchors act as attractors; message passing contracts the synthetic representation toward the class-$c$ manifold, mitigating residual generation noise.
Theorem 3.9 ( No confidence $\Rightarrow$ isolation ).
[…]
Interpretation. When agreement is uniformly low, the design quarantines the sample, preserving stability of the original graph.
4. Empirical Results
In this section, we present extensive experiments to validate our proposed method, emphasizing its benefits from both Semantic and Structural perspectives. We further evaluate the integrated framework to demonstrate its effectiveness in addressing long-tailed node classification. Specifically, our experiments aim to answer the following three research questions (RQs):
@@ 4. Empirical Results
4.1. Experimental Setup
4.1.1. Datasets
@@
Following prior studies on long-tailed graph learning (Fu et al., 2023; Zhou and Gong, 2023; Wang et al., 2021a), we use Cora (McCallum et al., 2000), PubMed (Sen et al., 2008), and Citeseer (Sen et al., 2008), along with three larger Amazon datasets, namely Photo, Computer, and Children (Yan et al., 2023), for scalability analysis.
To realistically simulate long-tailed distribution scenarios typical in real-world graphs, we follow the standard protocol from prior work (Yun et al., 2022; Wang et al., 2024; Fu et al., 2023; Zhou and Gong, 2023; Wang et al., 2021a; Zhao et al., 2021). Specifically, we randomly select 20 nodes from each majority (head) class, and $20\times\textit{imbalance\_ratio}$ nodes from each minority class. A summary of the detailed statistics of these datasets is provided in Table 9 from Appendix C.
4.1.2. Metrics
Following prior work on long-tailed graph classification (Yun et al., 2022; Wang et al., 2024; Fu et al., 2023; Zhou and Gong, 2023; Wang et al., 2021a; Zhao et al., 2021), we evaluate model performance using four standard metrics: balanced accuracy (bAcc), macro-F1 score (F1), geometric mean (GMean), and traditional accuracy (Acc). All experiments are repeated five times with different fixed seeds, and we report the mean performance across these runs.
To further assess boundary-enriching effects, we report the Boundary Coverage Ratio (BCR), as defined in Definition 3.4, and the Boundary Proximity Score (BPS). BPS is computed as the reciprocal of the Trust Score (Jiang et al., 2018), measuring the ratio between the distance to the nearest out-of-class centroid and the distance to the in-class centroid. Additionally, we report the In-Class Rate (ICR), defined as the percentage of samples predicted with the same label as their ground truth. Implementation details for these metrics are provided in Section 4.2.
4.1.3. Baselines
We compare our method against two categories of baselines (with details found in Appendix B):
  • LLM-based data augmentation: zero shot and few shot (Li et al., 2023; Yu et al., 2024)
  • Long-tailed graph learning: Oversampling (Chawla, 2003), SMOTE (Chawla et al., 2002),
    Embed-SMOTE(Ando and Huang, 2017), MixupForGraph (Wang et al., 2021a), GraphSMOTET(Zhao et al., 2021), GraphSMOTEO(Zhao et al., 2021), LTE4G(Yun et al., 2022), and HierTail (Wang et al., 2024).
To comprehensively evaluate our method, we design and implement the following three variants of SaVe-TAG, each corresponding to a specific VRM-based strategy. Detailed prompting strategies for each variant are provided in the Appendix B.
  • SaVe-TAGS: Inspired by SMOTE (Chawla et al., 2002), interpolates with nearest neighbors within the same class.
  • SaVe-TAGM: Extends SMOTE (Chawla et al., 2002) by interpolating embeddings with nearest neighbors from potentially different classes, inspired by Mixup techniques.
4.1.4. Implementation Details
4.2. Semantic-aware Benefits (RQ1)
On-Manifold.
To compare on-manifold LLM-based and off-manifold numeric interpolation, we evaluate node classification on multiple benchmarks. As shown in Figure 3, Embed leverages LLaMA 3.2-1B embeddings for interpolation, while SimCSE and SBERT serve as alternatives; our LLM-based method using LLaMA 3.2-1B consistently outperforms all numeric interpolation baselines.
Figure 3. Node classification performance (F1) under various implementation of SaVe-TAGS. Embed uses Llama3.2-1B embeddings for interpolation. SimCSE and SBERT are compared as alternative embedding baselines. Our LLM-based method with Llama3.2-1B consistently outperforms all the embedding-based baselines.
Boundary-enriching Effects.
We first visualize the embeddings from LLM-based and numeric interpolation. Figure 4 shows the UMAP projection of all six datasets, with the boundary of the original data circled. While numeric interpolated samples tend to stay within the original boundary, LLM-generated samples often extend beyond it, enlarging the VRM decision region.
We continue the evaluation with BCR and BPS metrics, as shown in the Figure 5. These metrics are computed for three settings: the original data, the combined set of original and interpolated data (Interp+Orig), and the interpolated data alone (Interp-Only). Across all benchmarks, our LLM-based interpolation consistently outperforms numeric interpolation on both metrics, suggesting that LLM-generated samples align more effectively with decision boundaries. Particularly in some cases, LLM-based interpolation even exceeds the original data in the BCR, highlighting its potential to extend and enrich the original decision boundary.
Figure 4. UMAP projection of the original and interpolated samples across various benchmarks—the original data boundary is outlined, with numeric samples mostly staying within and LLM-generated samples extending beyond.
Figure 5. Boundary Coverage Rate (BCR) and Boundary Proximity Score (BPS) for numeric (blue) vs. LLM (yellow) interpolation in Interp+Orig and Interp-Only settings across six datasets; the red dashed line marks the original baseline.
Figure 6. Average In-Class Rate (ICR) of augmented samples classified by an MLP trained on balanced data. Numerical interpolation achieves near-perfect class consistency (100% ICR), while LLM-based samples yield lower ICRs, reflecting increased proximity to decision boundaries.
In addition, we train a separate MLP classifier for each dataset to evaluate the class consistency of the augmented data. Each classifier is trained on a balanced dataset, where all classes have an equal number of training samples. The average In-Class Rate (ICR)—the proportion of augmented samples classified into their intended class—is reported in Figure 6. As we can see, Numerical interpolation achieves 100% ICR on Cora, PubMed, and Citeseer, and nearly 100% on others, reflecting high class consistency. In contrast, LLM-based samples show lower ICRs, indicating proximity to decision boundaries that causes occasional misclassification by MLP.
LLM as Vicinal Risk Minimization.
We evaluate the effectiveness of LLM-based interpolation within the Vicinal Risk Minimization (VRM) framework. As shown in Table 1, interpolation-based prompting significantly enhances the quality of synthesized textual data, resulting in improved performance on downstream node classification tasks.
Table 1. Comparison of VRM-based LLM interpolation with standard LLM augmentation. Subscripts O, M, and S denote the prompting strategies of their corresponding SaVe-TAG variants, but without edge assignment. Best and second-best results are emboldened and underlined, respectively.
@@ 4.2. Semantic-aware Benefits (RQ1)
Figure 7. Impact of Number of Assigned Edges. Left: The number of isolated nodes (log scale) decreases as the average number of assigned edges increases for the Photo, Computer, and Children datasets. Right: Effect of average assigned edges (log scale) on node classification for (a) Photo, (b) Computer, and (c) Children; performance is reported as changes from the baseline (n=1n{=}1).
4.3. Necessity of Confidence-based Edge Assignment (RQ2)